a designated hitter gets 8 hits in 9 games. if she continues hitting at this rate, how many hits will he get in 162 games?
Answer by macston(5194)

The hitter is hitting 8 hits/9 games or 8/9 hits per game.
8/9 hits/game x 162 games = 1296/9 hits = 144 hits